log(3x-5)-log(5x)=2
The key is no solution

Answer by scott8148(6628)

(3x - 5) / (5x) = 100

3x - 5 = 500x

-5 = 497x

-5/497 = x

logarithms are not defined for negative arguments , so the solution key is correct